# EvRT-DETR: Latent Space Adaptation of Image Detectors for Event-based Vision

### v2

@@ 4.3.3 Parameter Efficiency in Adaptation
Inspired by parameter-efficient techniques like LoRA [17], we investigate the minimal temporal module capacity needed for effective adaptation. Table 4 shows EvRT-DETR-B performance with varying hidden dimensions in the ConvLSTM modules.
Remarkably, reducing hidden dimensions to 64 features (a 4x reduction from our baseline) results in only modest performance degradation ($-0.6\%$ mAP), while requiring just 2.3 M additional parameters—merely $5.4\%$ of the base RT-DETR-B model’s 42.8M parameters. This configuration closely resembles LoRA-style adaptation, achieving substantial performance gains with minimal parameter overhead. Even our most parameter-efficient adaptation (M=64) outperforms all previous state-of-the-art approaches, demonstrating that effective temporal adaptation does not require extensive architectural modifications. Further increasing the hidden dimension to 512 features (42.9 M parameters, nearly doubling the model size) yields only a marginal $0.2\%$ mAP improvement indicating diminishing returns.
@@ 5 Conclusions
This work introduces I2EvDet, a framework that efficiently transforms image-based detectors into video-capable models through minimal modifications. Applied to event-based cameras, our approach demonstrates that mainstream object detection architectures can be effectively adapted to temporal domains without specialized redesign. RT-DETR trained on simple event representations achieves performance comparable to specialized methods, while our I2EvDet adapted model EvRT-DETR achieves state-of-the-art results on Gen1 ($+2.3$ mAP) and 1Mpx ($+1.4$ mAP) benchmarks. Our ablation studies provide insights on optimal temporal module design and parameter efficiency. These results suggest that bridging mainstream computer vision and specialized domains can be achieved through targeted adaptation rather than complete architectural redesign, potentially accelerating progress across diverse temporal visual domains.
@@ Acknowledgments
This manuscript has been authored by Brookhaven National Laboratory operated by Brookhaven Science Associates, LLC, under Contract Number DE-SC0012704 with the U.S. Department of Energy. The work was funded by the U.S. Department of Energy, National Nuclear Security Administration, Office of Defense Nuclear Nonproliferation Research and Development.
@@ Appendix A Training Setup
This section details the implementation of our I2EvDet framework as applied to event-based object detection. The training methodology reflects our two-stage adaptation approach: first establishing a strong spatial detector foundation, then incorporating minimal temporal processing while preserving the pre-trained model’s representation capabilities. This implementation demonstrates how mainstream detectors can be efficiently adapted to temporal data with minimal architectural changes
@@ A.1 RT-DETR
Our RT-DETR model is based on the reference PyTorch implementation111commit 5b628eaa0a2fc25bdafec7e6148d5296b144af85 [35, 36]. We explore two RT-DETR configurations: RT-DETR-T (corresponds to RT-DETR-ResNet18) and RT-DETR-B (corresponds to RT-DETR-ResNet50). Our RT-DETR configurations match the reference implementations, except two modifications: (1) we change the number of input channels of the backbones from 3 to 20, and (2) we do not use ImageNet pre-trained backbones. Table 5 summarizes the differences between the RT-DETR-ResNet18 and RT-DETR-ResNet50 configurations.
@@
Encoder Expansion ×$\times 0.5$$\times 1.0$
@@
Table 5: Comparison of the RT-DETR-T and RT-DETR-B Configurations
We train all RT-DETR models for $400{,}000$ iterations using the Adam optimizer [18] with a batch size of 32 and a learning rate of $2\times 10^{-4}$. Similar to the reference RT-DETR training, we maintain an exponential moving average (EMA) of the model weights with a momentum of $0.9999$. Unlike the reference RT-DETR training, we do not use any learning rate schedules, nor do we reduce the learning rate of the backbone relative to the encoder-decoder parts.
@@ A.2 EvRT-DETR
The EvRT-DETR-T and EvRT-DETR-B models extend RT-DETR-T and RT-DETR-B respectively by adding ConvLSTM modules [31]. Our ConvLSTM blocks use a hidden dimension of 256 (matching the encoder feature maps) and a kernel size of 3. The temporal module outputs are integrated into the base model through residual connections with a scaling factor of 1.0, as we found learnable scaling parameters (similar to ReZero [1]) provided no measurable performance benefit on the Gen1 dataset.
The ConvLSTM modules are trained jointly for $200{,}000$ iterations with Adam optimizer and a batch size of 8, while keeping the baseline RT-DETR models frozen. Each batch contains 8 short clips of consecutive frames: 21 frames for Gen1 dataset and 10 frames for 1Mpx dataset. The number of frames per clip follows RVT’s approach [11], except we increase the number of frames for the 1Mpx dataset from 5 to 10 for better performance. Each batch contains 4 randomly sampled clips and 4 consecutive clips from videos.
Unlike the baseline RT-DETR training with a constant learning rate and EMA averaging, we do not apply EMA averaging to the ConvLSTM modules and use a learning rate scheduler instead. For simplicity, we rely on the RVT-inspired one-cycle LR scheduler [32]. Specifically, we use PyTorch’s OneCycleLR implementation of the scheduler, with the following parameters: maximum LR $2\times 10^{-4}$, initial div_factor 20, final div_factor 500, pct_start 0.0005, annealing strategy ”linear”.
The ConvLSTM modules are trained in the same environment as the baseline RT-DETR networks, using a single NVIDIA RTX A6000 GPU.
A.3 Data Augmentation Strategy
@@
Our augmentation strategy employs a standard chain of geometric transformations and random erasing as summarized in Table 6. This approach bridges conventional image augmentation practices with the unique characteristics of event data, supporting our broader goal of adapting mainstream techniques to event-based vision.
@@
Random Rotation ±$\pm 30^{\circ}$ 0.6
Random Translation ±$\pm 0.5$ 0.6
@@
Random Shear ±$\pm 30^{\circ}$ 0.6
@@
Table 6 shows our augmentation strategy, which leverages standard transformations from the torchvision package (0.17.2[21]. Each augmentation is applied independently with its corresponding probability, and when applied, its magnitude is randomly sampled from the specified range. For Random Erasure, we preserve the object labels while erasing image regions, maintaining detection supervision even in partially occluded scenarios.
@@
For the temporal adaptation stage, we carefully preserve temporal consistency by applying identical geometric transformations (flip, rotation, scale, translation, shear) across all frames in a video clip, while still allowing per-frame variations through random erasing. This ensures that the temporal module learns meaningful motion patterns rather than artificially induced movements from inconsistent augmentations.
Appendix B Analysis of Base Detector Performance
This section explores key factors that influence the effectiveness of mainstream object detectors on event data, including data augmentation strategies and comparisons across different detector architectures.
B.1 Impact of Augmentation on Base Detector Performance.
@@ Appendix B Analysis of Base Detector Performance
B.2 Base Architecture Comparison for Event Data.
To investigate whether our findings about mainstream object detectors on event data extend beyond transformer-based models, we evaluated another popular architecture family - YOLOX [9], a CNN-based detector that has been used in other EBC applications [40, 11].
@@
RT-DETR-B (ours) $\mathbf{47.6}$ $\mathbf{76.3}$ $\mathbf{49.5}$
@@
Table 8: Performance comparison of different base detector architectures on individual event frames from the Gen1 dataset.
We trained YOLOX models on individual event frames using identical settings to our RT-DETR experiments. As shown in Table 8, YOLOX models achieve respectable performance on the Gen1 dataset, with YOLOX-X reaching 43.4 mAP, which remains below RT-DETR-B’s 47.6 mAP.
These results demonstrate that transformer-based architectures like RT-DETR provide significant advantages for this task compared to CNN-based approaches like YOLOX. The performance differential suggests that our I2EvDet framework benefits from starting with a transformer-based foundation, though the principle of adapting mainstream detectors to event data appears broadly applicable. Testing with the most recent YOLO iterations could be an interesting direction for future work, but was beyond the scope of our current study.
Appendix C Adaptation Considerations for Higher Resolution Event Data (1Mpx)
@@
C.1 Resolution Adaptation Strategy
@@
As shown in Table 9, nearest-neighbor interpolation substantially degrades RT-DETR performance, while bilinear interpolation yields optimal results. Interestingly, bicubic interpolation, despite its theoretical advantages for natural images, proves less effective for event data. These findings align with observations from AEC [23] and highlight the importance of selecting appropriate domain transfer techniques when adapting mainstream vision models to event data.
C.2 Temporal Context Length for Effective Adaptation
The temporal dimension represents a critical aspect of our adaptation framework. While prior work like RVT [11] used 5-frame clips for temporal training, our experiments indicated that expanded temporal context benefits the adaptation process.
@@ C.2 Temporal Context Length for Effective Adaptation
Table 10 demonstrates the improvement achieved by extending the temporal window to 10 frames during the adaptation phase. This finding suggests that providing longer temporal context during training allows the model to better capture persistent object representations across the temporal dimension, particularly important for event data where objects may be incompletely represented in shorter time windows.

### v3

@@ 4.4.2 Parameter Efficiency in Adaptation
Inspired by parameter-efficient techniques such as LoRA (Low-Rank Adaptation) [18], we investigate the minimal temporal module capacity needed for effective adaptation. Table 4 shows EvRT-DETR-B performance with varying hidden dimensions in the ConvLSTM modules.
Remarkably, reducing hidden dimensions to 64 features (a 4x reduction from our baseline) results in only modest performance degradation ($0.6\%$ mAP) while requiring just 2.3 M additional parameters – merely $5.4\%$ of the base RT-DETR-B model’s 42.8 M parameters. This configuration closely resembles LoRA-style adaptation, achieving substantial performance gains with minimal parameter overhead. Even our most parameter-efficient adaptation (M=64) outperforms all previous state-of-the-art approaches, demonstrating that effective temporal adaptation does not require extensive architectural modifications. Increasing the hidden dimension to 512 features (42.9 M parameters, nearly doubling the model size) yields only a marginal $0.2\%$ mAP improvement, indicating diminishing returns.
@@ 5 Conclusions
This work introduces I2EvDet, a framework that efficiently transforms image-based detectors into video-capable models using minimal modifications. Applied to event-based cameras, our approach demonstrates that mainstream object detection architectures can be effectively adapted to temporal domains without specialized redesign. RT-DETR trained on simple event representations achieves performance comparable to specialized methods, while our I2EvDet-adapted model EvRT-DETR achieves state-of-the-art results on Gen1 (+2.3+2.3 mAP) and 1Mpx (+1.4+1.4 mAP) benchmarks. Ablation studies provide insights about optimal temporal module design and parameter efficiency. These results suggest that bridging mainstream computer vision and specialized domains can be achieved through targeted adaptation rather than complete architectural redesign, potentially accelerating progress across diverse temporal visual domains.
@@ Acknowledgments
The authors extend their sincere gratitude to Yuewei Lin (Brookhaven Lab) for his valuable feedback and suggestions. This work presented in this paper was funded by the U.S. Department of Energy (DOE), National Nuclear Security Administration, Office of Defense Nuclear Nonproliferation Research and Development. The manuscript has been authored by Brookhaven National Laboratory managed by Brookhaven Science Associates, LLC under the Contract No. DE-SC0012704 with the U.S. Department of Energy.
@@ Appendix A Training Setup
This section details the implementation of the I2EvDet framework as applied to event-based object detection. The training methodology reflects a two-stage adaptation approach: first establishing a robust spatial detector foundation then incorporating minimal temporal processing while preserving the pre-trained model’s representation capabilities. This implementation demonstrates how mainstream detectors can be efficiently adapted to temporal data with minimal architectural changes
@@ A.1 RT-DETR
Our RT-DETR model is based on the reference PyTorch implementation111commit 5b628eaa0a2fc25bdafec7e6148d5296b144af85 [37, 38]. We explore two RT-DETR configurations: RT-DETR-T (corresponds to RT-DETR-ResNet18) and RT-DETR-B (corresponds to RT-DETR-ResNet50). The RT-DETR configurations match the reference implementations, except two modifications: (1) we change the number of input channels of the backbones from 3 to 20, and (2) we do not use ImageNet pre-trained backbones. Table 5 summarizes the differences between the RT-DETR-ResNet18 and RT-DETR-ResNet50 configurations.
@@
Encoder Expansion $\times 0.5$ $\times 1.0$
@@
Table 5: Comparison of the RT-DETR-T and RT-DETR-B Configurations.
We train all RT-DETR models for 400,000400{,}000 iterations using the Adam optimizer [19] with a batch size of 32 and a learning rate of $2\times 10^{-4}$. Similar to the reference RT-DETR training, we maintain an exponential moving average (EMA) of the model weights with a momentum of $0.9999$. Unlike the reference RT-DETR training, we do not use any learning rate schedules or reduce the learning rate of the backbone relative to the encoder-decoder parts.
@@ A.2 EvRT-DETR
The EvRT-DETR-T and EvRT-DETR-B models respectively extend RT-DETR-T and RT-DETR-B by adding ConvLSTM modules [33]. Our ConvLSTM blocks use a hidden dimension of 256 (matching the encoder feature maps) and a kernel size of 3. The temporal module outputs are integrated into the base model through residual connections with a scaling factor of 1.0 as we found learnable scaling parameters (similar to ReZero [1]) provide no measurable performance benefit on the Gen1 dataset.
The ConvLSTM modules are trained jointly for 200,000200{,}000 iterations with the Adam optimizer and a batch size of 8 while keeping the baseline RT-DETR models frozen. Each batch contains 8 short clips of consecutive frames: 21 frames for Gen1 dataset and 10 frames for 1Mpx dataset. The number of frames per clip follows RVT’s approach [12], except we increase the number of frames for the 1Mpx dataset from 5 to 10 for better performance. Each batch contains 4 randomly sampled clips and 4 consecutive clips from videos.
Unlike the baseline RT-DETR training with a constant learning rate and EMA averaging, we do not apply EMA averaging to the ConvLSTM modules and use a learning rate scheduler instead. For simplicity, we rely on the RVT-inspired one-cycle LR scheduler [34]. Specifically, we use PyTorch’s OneCycleLR implementation of the scheduler, with the following parameters: maximum LR $2\times 10^{-4}$, initial div_factor 20, final div_factor 500, pct_start 0.0005, and annealing strategy “linear.”
The ConvLSTM modules are trained in the same environment as the baseline RT-DETR networks using a single NVIDIA RTX A6000 GPU.
A.3 YOLOX Baselines
We implement YOLOX baselines to demonstrate the generalizability of our I2EvDet framework beyond RT-DETR. We use the reference YOLOX backbone and head implementations222commit ac58e0a5e68e57454b7b9ac822aced493b553c53 [9, 10]. The only architectural modification is changing the number of input channels of the backbone from 3 to 20.
To ensure fair comparison across architectures, we train YOLOX baselines using exactly the same setup as RT-DETR (cf. subsection A.1). Specifically, we train for 400,000400{,}000 iterations using the Adam optimizer with a batch size of 32 and a learning rate of $2\times 10^{-4}$.
The inference is performed with NMS threshold of $0.65$ and confidence threshold of $0.01$.
A.4 I2EvDet on YOLOX
Following our I2EvDet framework, we extend the YOLOX baseline with temporal processing capabilities. We freeze the pre-trained YOLOX parameters and insert ConvLSTM modules between the PAFPN [22] neck and detection head.
The ConvLSTM modules follow the same configuration as EvRT-DETR: hidden dimension of 256, kernel size of 3, and residual integration with scaling factor 1.0. Training follows the identical protocol as EvRT-DETR with 200,000200{,}000 iterations using the Adam optimizer and OneCycleLR scheduler (max LR $2\times 10^{-4}$). Each batch contains 4 random and 4 consecutive clips of 21 frames (Gen1) or 10 frames (1Mpx).
A.5 Notes on Two-Stage versus End-to-End Training
While developing the I2EvDet framework, we experimented with both end-to-end and two-stage training approaches. For the end-to-end training, we combined the RT-DETR model with the ConvLSTM temporal adapters and trained them jointly. Our experiments indicate that the two-stage training approach enables significantly faster convergence compared to end-to-end training with much higher stability. Moreover, in our experiments, the two-stage training has no performance downsides compared to the end-to-end training.
Based on these observations, we adopt the two-stage training approach in this work. An additional benefit of two-stage training is it enables experimentation with different temporal modules on the same spatial detector baseline, accelerating development and providing clear separation between spatial and temporal performance components.
We note that our end-to-end training experiments did not involve extensive hyperparameter exploration, so it is possible that alternative hyperparameter configurations may improve end-to-end training performance. However, the two-stage approach has proven more robust and practical for our framework development.
A.6 Data Augmentation Strategy
@@
Our augmentation strategy employs a standard chain of geometric transformations and random erasing as summarized in Table 6. This approach bridges conventional image augmentation practices with the unique characteristics of event data, supporting the broader goal of adapting mainstream techniques to event-based vision.
@@
Random Rotation $\pm 30^{\circ}$ 0.6
Random Translation $\pm 0.5$ 0.6
@@
Random Shear $\pm 30^{\circ}$ 0.6
@@
Table 6 shows our augmentation strategy, which leverages standard transformations from the torchvision package (0.17.2[23]. Each augmentation is applied independently with its corresponding probability. When applied, its magnitude is randomly sampled from the specified range. For Random Erasure, we preserve the object labels while erasing image regions, maintaining detection supervision even in partially occluded scenarios.
@@
For the temporal adaptation stage, we carefully preserve temporal consistency by applying identical geometric transformations (flip, rotation, scale, translation, shear) across all frames in a video clip while still allowing per-frame variations through random erasing. This ensures that the temporal module learns meaningful motion patterns rather than artificially induced movements from inconsistent augmentations.
Appendix B Supplementary Ablation Studies
This section provides additional ablation studies examining data augmentation strategies for event-based object detection and spatial context size in our temporal adaptation modules.
B.1 Impact of Augmentation on Base Detector Performance
@@
Appendix C Spatial Context Size in Temporal Processing
@@
EvRT-DETR-B (KS=1) $52.3$ $81.4$ $55.4$
EvRT-DETR-B (KS=3) $\mathbf{52.7}$ $\mathbf{82.0}$ $\mathbf{56.0}$
EvRT-DETR-B (KS=5) $52.0$ $81.3$ $54.8$
Table 8: Effect of ConvLSTM Kernel Size. Ablation study on Gen1 dataset showing optimal temporal adaptation performance with $3\times 3$ kernels.
Previous work on event-based object detection with ConvLSTM models, specifically RVT [12], found optimal performance with $1\times 1$ ConvLSTM kernels (effectively pointwise LSTMs). However, our adaptation approach shows different optimal characteristics. As shown in Table 8, EvRT-DETR achieves best performance with $3\times 3$ kernels, suggesting that spatial context is valuable when adapting frozen RT-DETR features. This demonstrates how adaptation design choices may differ from specialized architectures built from the ground up. Increasing to $5\times 5$ kernels degrades performance, suggesting that the additional parameters and receptive field expansion do not provide beneficial information for the task given the available training data.
Appendix D Detailed YOLOX Results
To demonstrate that our findings extend beyond transformer-based architectures, we evaluate our I2EvDet framework on YOLOX [9], a CNN-based detector that has been used in other EBC applications [42, 12].
Model mAP (%) $\text{mAP}_{50}$ (%) $\text{mAP}_{75}$ (%)
RT-DETR-B $\mathbf{47.6}$ $\mathbf{76.3}$ $\mathbf{49.5}$
@@ Appendix D Detailed YOLOX Results
EvRT-DETR-B $\mathbf{52.7}$ $\mathbf{82.0}$ $\mathbf{56.0}$
EvYOLOX-T 42.4 71.9 42.8
EvYOLOX-S 43.6 73.1 44.4
EvYOLOX-L 46.6 75.4 48.2
EvYOLOX-X 47.8 75.7 50.0
Table 9: Performance comparison of RT-DETR and YOLOX variants with and without our I2EvDet temporal adaptation on the Gen1 dataset. The consistent improvements across all architectures and model sizes demonstrate the generalizability of our adaptation framework beyond transformer-based detectors to CNN-based architectures.
We apply our two-stage training approach to multiple YOLOX variants, using identical training configurations to our RT-DETR experiments. Table 9 shows that YOLOX models achieve respectable baseline performance on the Gen1 dataset with YOLOX-X reaching 43.4 mAP, which remains below RT-DETR-B’s 47.6 mAP.
The I2EvDet framework provides substantial improvements across all YOLOX variants with gains ranging from 4.4 to 6.4 mAP. These consistent improvements demonstrate that temporal adaptation benefits extend across different architectural paradigms, validating our framework’s generalizability. While transformer-based RT-DETR achieves higher absolute performance than CNN-based YOLOX variants, both architecture families benefit significantly from our temporal adaptation approach.
These results confirm that the I2EvDet framework represents a general strategy for adapting image-based detectors to temporal domains with broad applicability beyond transformer architectures. Testing with the most recent YOLO iterations could be an interesting direction for future work but is beyond the scope of this current study.
Figure 5: Qualitative detection results on diverse 1Mpx automotive scenarios. Top row: RT-DETR predictions. Middle row: EvRT-DETR predictions. Bottom row: Ground truth annotations. Bounding box colors indicate object classes: red (cars), blue (pedestrians), orange (two-wheelers). Both methods perform well on dynamic scenes with moving objects, demonstrating the effectiveness of our base RT-DETR adaptation to event-based data across varied lighting conditions and object configurations
Appendix E Adaptation Considerations for Higher Resolution Event Data (1Mpx)
@@
E.1 Resolution Adaptation Strategy
@@
Table 10 depicts how the nearest-neighbor interpolation substantially degrades RT-DETR performance, while bilinear interpolation yields optimal results. Interestingly, despite its theoretical advantages for natural images, bicubic interpolation proves less effective for event data. These findings align with observations from AEC [25] and highlight the importance of selecting appropriate domain transfer techniques when adapting mainstream vision models to event data.
E.2 Temporal Context Length for Effective Adaptation
The temporal dimension represents a critical aspect of our adaptation framework. While prior work like RVT [12] uses 5-frame clips for temporal training, our experiments indicate that expanded temporal context benefits the adaptation process.
@@ E.2 Temporal Context Length for Effective Adaptation
Table 11 demonstrates the improvement achieved by extending the temporal window to 10 frames during the adaptation phase. This finding implies that providing longer temporal context during training allows the model to better capture persistent object representations across the temporal dimension, which is particularly important for event data where objects may be incompletely represented in shorter time windows.
Appendix F Comprehensive Evaluation Metrics
To provide complete experimental validation, this section presents detailed performance metrics and additional visualizations of our models across both Gen1 and 1Mpx datasets.
F.1 Detailed COCO Metrics
Model Data mAP ($\%$) $\text{mAP}_{50}$ ($\%$) $\text{mAP}_{75}$ ($\%$) mAR ($\%$) $\text{mAR}_{50}$ ($\%$) $\text{mAR}_{75}$ ($\%$)
RT-DETR-T (ours) Gen1 46.0 74.4 47.2 38.3 53.2 38.3
RT-DETR-B (ours) Gen1 47.5 76.3 49.5 39.9 54.6 42.1
EvRT-DETR-T (ours) Gen1 52.3 81.4 55.2 44.2 59.2 51.8
EvRT-DETR-B (ours) Gen1 $\mathbf{52.7}$ $\mathbf{82.0}$ $\mathbf{56.0}$ $\mathbf{45.1}$ $\mathbf{59.3}$ $\mathbf{53.3}$
RT-DETR-T (ours) 1Mpx 44.1 74.2 44.3 33.5 50.7 52.6
RT-DETR-B (ours) 1Mpx 45.2 75.1 46.0 35.1 51.8 53.5
EvRT-DETR-T (ours) 1Mpx 49.9 $\mathbf{81.0}$ 51.6 38.6 55.5 $\mathbf{62.4}$
EvRT-DETR-B (ours) 1Mpx $\mathbf{50.1}$ 80.9 $\mathbf{52.1}$ $\mathbf{39.0}$ $\mathbf{55.6}$ 61.7
Table 12: Complete COCO Evaluation Metrics. Comprehensive performance comparison showing mAP and mAR metrics at IoU thresholds of 0.5:0.95 (default), 0.5, and 0.75 for all model variants on Gen1 and 1Mpx datasets. Our temporal adaptation consistently improves both precision and recall across all thresholds and datasets, validating the robustness of the I2EvDet framework.
Table 12 provides comprehensive COCO evaluation metrics including mean Average Precision (mAP) and mean Average Recall (mAR) at different Intersection over Union (IoU) thresholds. I2EvDet’s temporal adaptation consistently improves performance across all metrics and datasets, demonstrating the robustness of our approach.
F.2 Additional Visualizations
Figure 5 presents qualitative detection results on diverse 1Mpx automotive scenarios across varied lighting conditions and object configurations. Both RT-DETR and EvRT-DETR demonstrate effective adaptation to event-based data representations, with comparable performance on dynamic scenes where objects generate sufficient event data through motion.
Figure 6: Temporal sequence demonstrating event-based detection challenges during motion transitions. A vehicle approaches an intersection, stops (creating sparse event data), then resumes motion. Frames shown every 100 frames for visualization clarity. Top row: RT-DETR predictions degrade significantly during stationary periods due to minimal event generation. Middle row: EvRT-DETR maintains more consistent detection by leveraging temporal memory from previous frames when objects were actively generating events. Bottom row: Ground truth annotations. Bounding box colors indicate object classes: red (cars), blue (pedestrians), orange (two-wheelers). While our temporal module substantially improves detection consistency during low-activity periods, challenges remain for heavily occluded objects, highlighting opportunities for future work
Figure 6 illustrates the critical advantage of temporal memory during motion transitions. As vehicles stop at an intersection and event generation becomes sparse, frame-based detection degrades significantly while our temporal adaptation maintains consistent object localization by leveraging historical information. This sequence exemplifies the fundamental challenge that motivates our I2EvDet framework and demonstrates its effectiveness in real-world automotive scenarios.